Alfred buys an old scooter for PhP 14 700 and spends PhP 1 800 on its repairs. If



he sells the scooter for PhP 15 800, what is his percent gain or percent loss?

1
Expert's answer
2021-12-13T06:31:51-0500

Total spend: a = 14700 + 1800 = 16500

Selling price: b = 15800

Income: c = b - a = 15800 - 16500 = -700

Percent loss: "{\frac c a}*100={\frac {700} {16500}}*100={\frac {140} {33}}\approx4.24" %